#2 near the ground at the north magnetic pole, the earth’s magnetic field is directed downward vertically and has a str...

#2
near the ground at the north magnetic pole, the earth’s magnetic field is directed downward vertically and has a strength of 0.57 G. An electron at the north pole is projected from the ground into this magnetic field with a velocity of 1.4 km/s vertically upward and 1.9 km/s horizontally to your right.
A) find the magnitude of the force acting on this electron
B) find the electron’s acceleration (magnitude and direction)
Answer #1

- B -0.579 (-) ū - 14 komys +1.9 km/s 1 B q = - 1-6 x 10-la e groura A) IG= 10-4 T. - B = 0.57x10h T (*) u a 1.4 x 103 m/s Ŕ